OpenStudy (anonymous):

Hello! I need to simplify a product, but am not sure how! I will put the product in the answers. Please help! And don't just give me the answer, explain how to get it! Alright thanks! c:

OpenStudy (anonymous):

\[\left( 5\sqrt{2} -2\sqrt{3}\right)^2\] there is the product! c:

OpenStudy (anonymous):

@myininaya help please?

OpenStudy (anonymous):

expand the equation by removing the brackets

OpenStudy (anonymous):

would i just rewrite it without the brackets?

OpenStudy (anonymous):

(52√−23√)(52√−23√)

OpenStudy (anonymous):

just like that. dont mind the 2 dat was omited, it was a mistake

OpenStudy (anonymous):

then multiply using 'surds'

OpenStudy (anonymous):

okay it's fine haha. what is surds?

OpenStudy (anonymous):

i dont really know how to explain it. tis about multiplying or adding fractions with roots

OpenStudy (anonymous):

okay, so would I multiply the two 5's and 2's and come up with this...\[(25\sqrt{2}^2-4\sqrt{3}^2)^2\]

OpenStudy (anonymous):

nooo

OpenStudy (anonymous):

yup I've got no idea what I'm doing haha

OpenStudy (anonymous):

\[(5\sqrt{2} - 2\sqrt{3})(5\sqrt{2} - 2\sqrt{3})\]

OpenStudy (anonymous):

you will get:

OpenStudy (anonymous):

ohh... that's the answer isn't it? :p i'm a little slow

OpenStudy (anonymous):

\[10\sqrt{2} + (5\sqrt{2} * 2\sqrt{3}) - (2\sqrt{3} * 5\sqrt{2}) + 4\sqrt{3}\]

OpenStudy (anonymous):

you will then need to simplify

OpenStudy (anonymous):

hmm okay

OpenStudy (anonymous):

by opening up the brackets

OpenStudy (anonymous):

\[10\sqrt{2} - 10\sqrt{6} + 10\sqrt{6} + 4\sqrt{3}\]

OpenStudy (anonymous):

\[10\sqrt{6}\] cancels out

OpenStudy (anonymous):

the answer would be: \[10\sqrt{2} + 4\sqrt{3}\]

OpenStudy (anonymous):

Oh! That makes so much sense! thank you so much! (:

OpenStudy (anonymous):

you er welcome. just give me a medal :)